OpenStudy (zzzzzz):

x2-16x-192 split the middle term?

OpenStudy (anonymous):

(x + 8) • (x - 24) = 0

OpenStudy (anonymous):

(x + 8) • (x - 24) = 0

OpenStudy (anonymous):

(x+8)(x-24)=0

OpenStudy (anonymous):

x=-8 and x=24

OpenStudy (zzzzzz):

Thanks a lot brother